# Collocation methods for nonlinear differential equations on low-rank manifolds

### v2

@@ Abstract
We introduce new methods for integrating nonlinear differential equations on low-rank manifolds. The methods are based on oblique projectors onto the tangent spaces of low-rank manifolds with an interpolation property that collocates the differential equation on the manifold. These projections yield evolution equations that enable low-rank time integration of any vectorfield that can be evaluated entry-wise. The proposed methods do not require the vectorfield to have a low-rank structure, thus overcoming significant limitations of dynamical low-rank approximation based on orthogonal projection. To construct the oblique projectors, we devise a sparse tensor sampling algorithm based on the discrete empirical interpolation method (DEIM) that parameterizes tensor train manifolds and their tangent spaces with cross interpolation. Using these projectors we propose two time integration schemes on low-rank tensor train manifolds. The first is based on a time-dependent tensor cross interpolation representation of the solution and the second is a direct generalization of the well-known orthogonal projector-splitting integrator to a oblique projectors. We demonstrate the proposed methods with applications to several tensor differential equations arising from the discretization of partial differential equations.
@@ 1 Introduction
Two classes of algorithms for integrating (2) on smooth tensor manifolds are step-truncation methods [42] and dynamical low-rank methods [10]. Step-truncation methods allow the solution rank to naturally increase in a controlled manner during a time step before truncating back to the desired rank. Dynamical low-rank methods integrate the solution on a fixed-rank manifold by projecting $G(X,t)$ onto a tangent space of the manifold at each time $t$. Both methods aim to compute the best approximate solution to (2) on a fixed-rank tensor manifold at each time $t$ and are consistent with each other as the temporal step-size approaches zero (see, e.g., [9, Section 3.3]). These methods have been utilized for several applications including uncertainty quantification [2, 45], plasma physics [54, 17], numerical approximation of functional differential equations [53, 44] and machine learning [48, 46], and substantial research efforts have recently been made to improve their accuracy, efficiency, and robustness. These efforts have resulted in several innovations including rank-adaptive integrators [9, 5, 42], implicit low-rank methods [43, 36, 50], conservative low-rank methods [25, 3, 18, 16] and coordinate-adaptive low-rank methods [11, 12].
@@
The rest of this paper is organized as follows. In Section 2 we briefly recall the TT format and orthogonalization of tensors in the TT format. In Section 3 we recall the orthogonal projector onto the TT tangent space and introduce new oblique projectors onto the tangent space. Then we describe a special class of oblique projectors with a cross interpolation property. In Section 4 we present a new index selection algorithm, referred to as TT-cross-DEIM, for constructing oblique projectors onto the TT tangent space. We show that indices obtained with the TT-cross-DEIM algorithm define well-conditioned oblique tangent space projectors and parameterize TT manifolds with cross interpolation. In Section 5 we introduce new dynamical low-rank time integration schemes for (2) using oblique tangent space projectors. In Section 6 we demonstrate the proposed dynamical low-rank methods and compare the results with existing time integration methods on low-rank tensor manifolds. The main findings are summarized in Section 7.
2 Tensor train format
To introduce notation and make our presentation self-contained we begin with a brief review of the tensor train (TT) format and orthogonal TT representations. For more details on the TT format we refer the reader to [38]. The $k$th unfolding of a tensor $Y\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ is the matrix $\bm{Y}^{\langle k\rangle}\in\mathbb{R}^{(n_{1}\cdots n_{k})\times(n_{k+1}\cdots n_{d})}$ obtained by reshaping the entries of $Y$, where the row and column indices form a colexicographical ordering. The TT-rank of $Y$ is defined as the vector $\bm{r}=(1,r_{1},\ldots,r_{d},1)$ where $r_{k}$ is the rank of the unfolding matrix $\bm{Y}^{\langle k\rangle}$. Any tensor $Y$ with TT-rank $\bm{r}$ can be represented in the TT format as
@@ 2 Tensor train format
where each $C_{k}$ is a $r_{k-1}\times n_{k}\times r_{k}$ tensor referred to as a TT-core and $C_{k}(i_{k})$ is a $r_{k-1}\times r_{k}$ matrix for a fixed index $i_{k}$. Each TT-core has a left unfolding matrix $\bm{C}_{k}^{\langle l\rangle}\in\mathbb{R}^{r_{k-1}n_{k}\times r_{k}}$ and a right unfolding matrix $\bm{C}_{k}^{\langle r\rangle}\in\mathbb{R}^{r_{k-1}\times n_{k}r_{k}}$ obtained by reshaping the elements of $C_{k}$
@@
The left and right unfoldings of each TT-core is full rank whenever $Y$ has TT-rank $\bm{r}$.
            (a)                                                                  (b)
To simplify notation of tensors in the TT format we often omit indices so that (3) is replaced by $Y=C_{1}C_{2}\cdots C_{d}$. To obtain even more compact representations of tensors in the TT format, we define partial products of TT-cores $C_{\leq k}\in\mathbb{R}^{n_{1}\times\cdots\times n_{k}\times r_{k}}$ and $C_{>k}\in\mathbb{R}^{r_{k}\times n_{k+1}\times\cdots\times n_{d}}$ with entries
@@
where $\bm{C}_{\leq k}\in\mathbb{R}^{(n_{1}\cdots n_{k})\times r_{k}}$ and $\bm{C}_{>k}\in\mathbb{R}^{(n_{k+1}\cdots n_{d})\times r_{k}}$ to write the $k$th unfolding matrix of $Y$ as $\bm{Y}^{\langle k\rangle}=\bm{C}_{\leq k}\bm{C}_{>k}^{\text{T}}$.

### v3

@@ Abstract
We introduce new methods for integrating nonlinear differential equations on low-rank manifolds. These methods rely on interpolatory projections onto the tangent space, enabling low-rank time integration of vector fields that can be evaluated entry-wise. A key advantage of our approach is that it does not require the vector field to exhibit low-rank structure, thereby overcoming significant limitations of traditional dynamical low-rank methods based on orthogonal projection. To construct the interpolatory projectors, we develop a sparse tensor sampling algorithm based on the discrete empirical interpolation method (DEIM) that parameterizes tensor train manifolds and their tangent spaces with cross interpolation. Using these projectors, we propose two time integration schemes on low-rank tensor train manifolds. The first scheme integrates the solution at selected interpolation indices and constructs the solution with cross interpolation. The second scheme generalizes the well-known orthogonal projector-splitting integrator to interpolatory projectors. We demonstrate the proposed methods with applications to several tensor differential equations arising from the discretization of partial differential equations.
@@ 1 Introduction
Two classes of algorithms for integrating (2) on smooth tensor manifolds are step-truncation methods [42] and dynamical low-rank methods [10]. Step-truncation methods allow the solution rank to naturally increase in a controlled manner during a time step before truncating back to the desired rank. Dynamical low-rank methods integrate the solution on a fixed-rank manifold by projecting $G(X,t)$ onto a tangent space of the manifold at each time $t$. Both methods aim to efficiently compute the best approximate solution to (2) on a fixed-rank tensor manifold at each time $t$ and are consistent with each other as the temporal step-size approaches zero (see, e.g., [9, Section 3.3]). These methods have been utilized for several applications including uncertainty quantification [2, 45], plasma physics [54, 17], numerical approximation of functional differential equations [53, 44] and machine learning [48, 46], and substantial research efforts have recently been made to improve their accuracy, efficiency, and robustness. These efforts have resulted in several innovations including rank-adaptive integrators [9, 5, 42], implicit low-rank methods [43, 36, 50], conservative low-rank methods [25, 3, 18, 16] and coordinate-adaptive low-rank methods [11, 12].
@@
The rest of this paper is organized as follows. In Section 2 we introduce interpolatory tangent space projectors and the proposed dynamical low-rank methods on matrix manifolds ($d=2$). In Section 3 we briefly recall the TT format and orthogonalization of tensors in the TT format. In Section 4 we recall the orthogonal projector onto the TT tangent space and introduce new oblique projectors onto the tangent space. Then we describe a special class of oblique projectors with a cross interpolation property. In Section 5 we present a new index selection algorithm, referred to as TT-cross-DEIM, for constructing oblique projectors onto the TT tangent space. We show that indices obtained with the TT-cross-DEIM algorithm define oblique tangent space projectors and parameterize TT manifolds with cross interpolation. In Section 6 we introduce new dynamical low-rank time integration schemes for (2) using oblique tangent space projectors. In Section 7 we demonstrate the proposed dynamical low-rank methods and compare the results with existing time integration methods on low-rank tensor manifolds. The main findings are summarized in Section 8.
2 Dynamical low-rank matrix approximation
Figure 1: A sketch of the low-rank manifold $\mathcal{M}_{\bm{r}}$ and its tangent space $T_{Y}\mathcal{M}_{\bm{r}}$ at the point $Y\in\mathcal{M}_{\bm{r}}$. Also depicted is $Z\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ and its orthogonal projection $\widehat{P}_{Y}Z$ and oblique projection $P_{Y}Z$ onto the tangent space $T_{Y}\mathcal{M}_{\bm{r}}$. The orthogonal projection is the best approximation of $Z$ on the tangent space with respect to the Frobenius norm but is impractical compute when $G$ lacks rank structure. The oblique projection is a quasi-optimal approximation of $Z$ on the tangent space that is efficient to compute for any $G$ that can be evaluated entry-wise. Such oblique projectors allow us to efficiently apply dynamical low-rank methods to a broad class of nonlinear differential equations.
Before introducing dynamical low-rank tensor approximation, we describe the proposed low-rank methods for matrices ($d=2$). The goal is to find an approximate solution $Y(t)$ to (2) that lies on the manifold $\mathcal{M}_{r}$ of rank-$r$ matrices for all time $t$. A rank-$r$ matrix can be expressed through left and right factor matrices with dimensions $n_{1}r$ and $n_{2}r$, respectively. Thus approximating the solution to (2) with $Y(t)\in\mathcal{M}_{r}$ reduces the storage cost from $n_{1}n_{2}$ to $(n_{1}+n_{2})r$. Dynamical low-rank methods integrate the approximate solution $Y(t)$ on the manifold $\mathcal{M}_{r}$ by projecting (2) onto a tangent space of $\mathcal{M}_{r}$ at each time $t$. Assuming the initial condition $X_{0}$ belongs to $\mathcal{M}_{r}$, this procedure yields the evolution equation
$\frac{dY(t)}{dt}=P_{Y(t)}G(Y(t),t),\qquad Y(0)=X_{0},$ (3)
where $P_{Y(t)}:\mathbb{R}^{n_{1}\times n_{2}}\to T_{Y(t)}\mathcal{M}_{r}$ projects onto the tangent space of $\mathcal{M}_{r}$ at $Y(t)$. The solution to (3) remains on $\mathcal{M}_{r}$ for all $t\geq 0$ and serves as an approximate solution to (2). Classical dynamical low-rank methods use an orthogonal tangent space projector to minimize the error of the approximation in the Frobenius norm at each time $t$. Let $U(t)$ and $V(t)$ be matrices whose columns form orthonormal bases for the range and co-range of $Y(t)$, respectively. Such matrices can be obtained, for example, from the SVD of $Y(t)$. Then the orthogonal tangent space projector can be expressed as [31]
$\widehat{P}_{Y}Z=ZVV^{\text{T}}-UU^{\text{T}}ZVV^{\text{T}}+UU^{\text{T}}Z,$ (4)
where $Z\in\mathbb{R}^{n_{1}\times n_{2}}$ and we suppressed dependence on $t$ for simplicity. When $Z=G(Y(t),t)$ computing the projection (5) at each time step to integrate (3) can be computationally expensive, especially if $G$ does not have low-rank structure. For most nonlinear functions $G$, computing such orthogonal projection has computational cost scaling as $n_{1}n_{2}$, making the integration of the approximate solution $Y(t)\in\mathcal{M}_{r}$ as expensive as solving for the full solution $X(t)$ using standard methods.
2.1 Interpolatory dynamical low-rank approximation
Algorithm 1 DEIM index selection (adapted from [49])
1:$\bm{V}\in\mathbb{R}^{n\times r}$ with $n\geq r$
2:$\bm{l}$, a vector with $r$ distinct indices from $\{1,\ldots,n\}$
3:$\bm{v}=\bm{V}(:,1)$
[…]
5:for $j=2,3,\ldots,r$  do
6:     $\bm{v}=\bm{V}(:,j)$
7:     $\bm{c}=\bm{V}(l,1:j-1)^{-1}\bm{v}(\bm{l})$
8:     $\bm{r}=\bm{v}-\bm{V}(:,1:j-1)\bm{c}$
9:     $[~{},l_{j}]=\max(|\bm{r}|)$
10:     $\bm{l}=[\bm{l};l_{j}]$
11:end for
To develop efficient dynamical low-rank integrators for nonlinear $G$ we propose a new interpolatory tangent space projector $P_{Y}$ to replace the orthogonal projector. This interpolatory projector is a specialized oblique tangent space projector, obtained by replacing the orthogonal projectors $UU^{\text{T}}$ and $VV^{\text{T}}$ in (4) with oblique projectors onto the same spaces. A general form of these oblique projectors are $U(A^{\text{T}}U)^{-1}A^{\text{T}}$ and $V(B^{\text{T}}V)^{-1}B^{\text{T}}$ where $A\in\mathbb{R}^{n_{1}\times r}$ and $B\in\mathbb{R}^{n_{2}\times r}$ are any matrices such that $(A^{\text{T}}U)$ and $(B^{\text{T}}V)$ are invertible. Interpolatory projectors onto the columns of $U$ and $V$ are obtained by selecting $A$ and $B$ as specific columns of the identity matrix with appropriate dimensions. Specifically $A=I_{n_{1}}(:,\mathcal{I})$ and $B=I_{n_{2}}(:,\mathcal{J})$ where $\mathcal{I}$ contains $r$ indices from $\{1,\ldots,n_{1}\}$ and $\mathcal{J}$ contains $r$ indices from $\{1,\ldots,n_{2}\}$. The matrix $A$ can extract $r$ rows from $Z$ via left multiplication $A^{\text{T}}Z=Z(\mathcal{I},:)$ and the matrix $B$ can extract $r$ columns from $Z$ via right multiplication $ZB=Z(:,\mathcal{J})$. Moreover, $A$ can be defined by choosing $r$ indices in the set $\mathcal{I}$, and similarly, $B$ can be defined by choosing $r$ indices in the set $\mathcal{J}$. Such indices are to be chosen so that $A^{\text{T}}U=U(\mathcal{I},:)$ and $B^{\text{T}}V=V(\mathcal{J},:)$ are invertible, and ideally with a small condition number. This can be achieved with a sparse sampling algorithm such as the discrete empirical interpolation method (DEIM). The DEIM, summarized in Algorithm 1, is a greedy algorithm that selects an index for each column of $U$ or $V$ to minimize the condition number of the interpolatory projector as much as possible. Other sparse sampling strategies, such as Q-DEIM or oversampling methods, can also be used and may yield better-conditioned interpolatory projectors than DEIM in certain cases. Replacing the orthogonal projectors in (4) with interpolatory projectors onto the columns of $U$ and $V$ results in an interpolatory projector onto the tangent space $T_{Y}\mathcal{M}_{r}$ of the form
$P_{Y}Z=Z(:,\mathcal{J})V(\mathcal{J},:)^{-\text{T}}V^{\text{T}}-UU(\mathcal{I},:)^{-1}Z(\mathcal{I},\mathcal{J})V(\mathcal{J},:)^{-\text{T}}V^{\text{T}}+UU(\mathcal{I},:)^{-1}Z(\mathcal{I},:).$ (5)
It is easy to verify that $(P_{Y}Z)(i,j)=Z(i,j)$ whenever $i\in\mathcal{I}$ or $j\in\mathcal{J}$. In other words, the projection $P_{Y}Z$ interpolates $Z$ along rows and columns specified by the index sets $\mathcal{I}$ and $\mathcal{J}$.
2.1.1 Matrix cross integrator
A straightforward low-rank time integration scheme can be derived by evaluating the dynamical low-rank evolution equation (3) at the sampled indices $\mathcal{I}$ and $\mathcal{J}$ and leveraging the interpolation property. This leads to the system of evolution equations
$\frac{dY(\mathcal{I}(t),:,t)}{dt}$ $=G_{Y}(\mathcal{I}(t),:,t),$ (6)
$\frac{dY(:,\mathcal{J}(t),t)}{dt}$ $=G_{Y}(:,\mathcal{J}(t),t),$
where $G_{Y}(t)=G(Y(t),t)$. Equation (6) consists of $(n_{1}+n_{2})r$ coupled nonlinear differential equations governing the evolution of a subset of the entries in the approximate solution, which can be integrated using standard explicit or implicit methods. The indices $\mathcal{I}(t)$ and $\mathcal{J}(t)$ that define the interpolatory projector (5) are chosen at each time $t$ to ensure that the projector remains well-defined during time integration. If $G$ arises from the spatial discretization of a PDE (1) involving differential operators, evaluating $G_{Y}\left(\mathcal{I},:,t\right)$ and $G_{Y}\left(:,\mathcal{J},t\right)$ requires entries of $Y$ at indices adjacent to $\mathcal{I}$ and $\mathcal{J}$. The values of $Y$ at adjacent indices can always be obtained by constructing the solution $Y(t)$ as a low-rank matrix using CUR decomposition. For example given $Y(\mathcal{I},:)$ and $Y(:,\mathcal{J})$ at any time $t$ the approximate solution $Y$ can be obtained as
$Y=Y(:,\mathcal{J})Y(\mathcal{I},\mathcal{J})^{-1}Y(\mathcal{I},:).$ (7)
If $Y=U\Sigma V^{\text{T}}$ is the SVD and the indices $\mathcal{I}$ and $\mathcal{J}$ defining the interpolatory projector are chosen so that $U(\mathcal{I},:)$ and $V(\mathcal{J},:)$ are well-conditioned, then constructing $Y$ using (7) involves $Y(\mathcal{I},\mathcal{J})^{-1}=V(\mathcal{J},:)^{-\text{T}}\Sigma^{-1}U(\mathcal{I},:)^{-1}$. Thus the condition number of the middle matrix is inversely proportional to the smallest singular value of $Y$. Instead of constructing $Y$ in this way, if we first take a QR decomposition $Y(:,\mathcal{J})=QR$, then we can write $Y(\mathcal{I},\mathcal{J})=Q(\mathcal{I},:)R$ and the CUR formula (7) becomes
$Y=QQ(\mathcal{I},:)^{-1}Y(\mathcal{I},:),$ (8)
which is an interpolatory projection of $Y(\mathcal{I},:)$ onto the orthonormal basis $Q$. In particular the stability of constructing the solution using (8) depends on the condition of the interpolatory projector and is independent of the singular values of $Y$. Computing the right hand side of (6), and hence integrating the system, is efficient for any nonlinear $G$ that can be evaluated entry-wise, regardless of its low-rank structure. This enables us to efficiently compute dynamical low-rank approximations for problems where orthogonal tangent space projection is too expensive.
2.1.2 Projector-splitting integrator
An alternative approach to the matrix-cross integrator presented above for integrating (3) is to apply a standard splitting method, similar to the integrators proposed for orthogonal projectors in [31]. Since (3) is a sum of three terms, applying Lie-Trotter splitting yields three substeps commonly referred to as K-, S-, and L-step. Beginning from the rank-$r$ decomposition of the approximate solution $Y(t_{0})=U(t_{0})S(t_{0})V^{\text{T}}(t_{0})$ at time $t_{0}$, each step updates a single factor matrix. After all three steps we obtain the low-rank factors for the approximate solution $Y(t_{1})=U(t_{1})S(t_{1})V^{\text{T}}(t_{1})$ at time $t_{1}=t_{0}+\Delta t$. The substeps for interpolatory projector-splitting integrator are as follows. We denote by $\texttt{DEIM}(\cdot)$ a subroutine that takes a matrix of size $n\times r$ as input and outputs a collection of $r$ indices computed with the DEIM index selection (Algorithm 1).
  • 1.
    K-step: update $U(t_{0})\to U(t_{1})$ and $S(t_{0})\to R(t_{1})$.
    Compute interpolation indices $\mathcal{J}=\texttt{DEIM}(V(t_{0}))$. Then integrate the $n_{1}\times r$ differential equation
    $\frac{dK(t)}{dt}=G_{K}\left(:,\mathcal{J},t\right)\left[V\left(\mathcal{J},:,t_{0}\right)\right]^{-\text{T}},\qquad K(t_{0})=U(t_{0})S(t_{0}),$ (9)
    where $G_{K}(t)=G\left(K(t)V(t_{0})^{\text{T}},t\right)$ from $t_{0}$ to $t_{1}$, and perform a QR-decomposition $K(t_{1})=U(t_{1})R(t_{1})$.
  • 2.
    S-step: update $R(t_{1})\to\tilde{S}(t_{1})$.
    Compute interpolation indices $\mathcal{I}=\texttt{DEIM}(U(t_{1}))$. Then integrate the $r\times r$ matrix differential equation
    $\frac{d\tilde{S}(t)}{dt}=-\left[U\left(\mathcal{I},:,t_{1}\right)\right]^{-1}G_{S}\left(\mathcal{I},\mathcal{J},t\right)\left[V\left(\mathcal{J},:,t_{0}\right)\right]^{-\text{T}},\qquad\tilde{S}(t_{0})=R(t_{1}),$ (10)
    where $G_{S}(t)=G\left(U(t_{1})S(t)V(t_{0})^{\text{T}},t\right)$ from $t_{0}$ to $t_{1}$.
  • 3.
    L-step: update $V(t_{0})\to V(t_{1})$ and $\tilde{S}(t_{1})\to S(t_{1})$.
    Integrate the $n_{2}\times r$ matrix differential equation
    $\frac{dL(t)}{dt}=G_{L}\left(\mathcal{I},:,t\right)^{\text{T}}\left[U\left(\mathcal{I},:,t_{1}\right)\right]^{-\text{T}},\qquad L(t_{0})=V(t_{0})\tilde{S}(t_{1})^{\text{T}},$ (11)
    where $G_{L}(t)=G\left(U(t_{1})L(t)^{\text{T}},t\right)$ from time $t_{0}$ to $t_{1}$, and perform a QR-decomposition $L(t_{1})=V(t_{1})S(t_{1})^{\text{T}}$.
Similar to the matrix-cross integrator, the interpolatory projector-splitting integrator requires evaluating the output of $G$ at only a subset of $nr$ indices in the K- and L-step and $r^{2}$ indices in the S-step. These evaluations can be performed efficiently for any $G$ that can be evaluated entry-wise. Meanwhile the corresponding steps of the orthogonal projector-splitting integrator (summarized in [6]) involve inner products involving the output of $G$ that are only efficient to compute when $G$ has low-rank structure. The difference between the interpolatory projector-splitting integrator and the matrix cross integrator is the order in which interpolatory projection and time integration are performed. The former applies the interpolatory projection on the vector-field and then integrates the factor matrices of the solution. The latter integrates the solution at the specified indices and then performs an interpolatory projection onto the updated basis $Q$ in (8).
3 Tensor train (TT) format
In the following sections we propose dynamical low-rank approximation with interpolatory projections for tensors $(d\geq 2)$ in the tensor train (TT) format. When $d=2$ the TT methods described hereafter reduce to the matrix methods discussed in Section 2. We begin with a brief review of the TT format and orthogonal TT representations. For a more detailed introduction to the TT format we refer the reader to [38]. Throughout the remainder of this paper, matrices are denoted by boldface letters, while tensors are denoted by regular (non-bold) letters. The $k$th unfolding of a tensor $Y\in\mathbb{R}^{n_{1}\times\cdots\times n_{d}}$ is the matrix $\bm{Y}^{\langle k\rangle}\in\mathbb{R}^{(n_{1}\cdots n_{k})\times(n_{k+1}\cdots n_{d})}$ with rows and columns indexed colexicographically. The TT-rank of $Y$ is defined as the vector $\bm{r}=(1,r_{1},\ldots,r_{d},1)$ where $r_{k}$ is the rank of the unfolding matrix $\bm{Y}^{\langle k\rangle}$. Any tensor $Y$ with TT-rank $\bm{r}$ can be represented in the TT format as
@@ 3 Tensor train (TT) format
where each $C_{k}$ is a $r_{k-1}\times n_{k}\times r_{k}$ tensor referred to as a TT-core and $C_{k}(i_{k})$ is a $r_{k-1}\times r_{k}$ matrix for a fixed index $i_{k}$. Each TT-core has a left unfolding matrix $\bm{C}_{k}^{\langle l\rangle}\in\mathbb{R}^{r_{k-1}n_{k}\times r_{k}}$ and a right unfolding $\bm{C}_{k}^{\langle r\rangle}\in\mathbb{R}^{r_{k-1}\times n_{k}r_{k}}$ obtained by reshaping the elements of $C_{k}$
@@
The left and right unfoldings of each TT-core is full rank whenever $Y$ has TT-rank $\bm{r}$. To simplify notation of tensors in the TT format we often omit indices so that (12) is replaced by $Y=C_{1}C_{2}\cdots C_{d}$. To obtain even more compact representations, we define partial products of TT-cores $C_{\leq k}\in\mathbb{R}^{n_{1}\times\cdots\times n_{k}\times r_{k}}$ and $C_{>k}\in\mathbb{R}^{r_{k}\times n_{k+1}\times\cdots\times n_{d}}$ with entries
@@
where $\bm{C}_{\leq k}\in\mathbb{R}^{(n_{1}\cdots n_{k})\times r_{k}}$ and $\bm{C}_{>k}\in\mathbb{R}^{(n_{k+1}\cdots n_{d})\times r_{k}}$, which allows us to write the $k$th unfolding matrix of $Y$ as $\bm{Y}^{\langle k\rangle}=\bm{C}_{\leq k}\bm{C}_{>k}^{\text{T}}$.
            (a)                                                                  (b)